### v1

# Benchmark of segmentation techniques for pelvic fracture in CT and X-ray: Summary of the PENGWIN 2024 challenge

Abstract
The segmentation of pelvic fracture fragments in CT and X-ray images is crucial for trauma diagnosis, surgical planning, and intraoperative guidance. However, accurately and efficiently delineating the bone fragments remains a significant challenge due to complex anatomy and imaging limitations. The PENGWIN challenge, organized as a MICCAI 2024 satellite event, aimed to advance automated fracture segmentation by benchmarking state-of-the-art algorithms on these complex tasks. A diverse dataset of 150 CT scans was collected from multiple clinical centers, and a large set of simulated X-ray images was generated using the DeepDRR method. Final submissions from 16 teams worldwide were evaluated under a rigorous multi-metric testing scheme. The top-performing CT algorithm achieved an average fragment-wise intersection over union (IoU) of 0.930, demonstrating satisfactory accuracy. However, in the X-ray task, the best algorithm attained an IoU of 0.774, highlighting the greater challenges posed by overlapping anatomical structures. Beyond the quantitative evaluation, the challenge revealed methodological diversity in algorithm design. Variations in instance representation, such as primary-secondary classification versus boundary-core separation, led to differing segmentation strategies. Despite promising results, the challenge also exposed inherent uncertainties in fragment definition, particularly in cases of incomplete fractures. These findings suggest that interactive segmentation approaches, integrating human decision-making with task-relevant information, may be essential for improving model reliability and clinical applicability.
keywords:
Pelvic fracture, Image segmentation, Deep learning, Challenge
journal: Medical Image Analysis
[…]
[Rossum]organization=Beijing Rossum Robot Technology Co., Ltd., city=Beijing, country=China
[…]
[Beihang]organization=Key Laboratory of Biomechanics and Mechanobiology, Ministry of Education, Beijing Advanced Innovation Center for Biomedical Engineering, School of Biological Science and Medical Engineering, Beihang University, city=Beijing, postcode=100083, country=China
[…]
[JHU]organization=Department of Computer Science, Johns Hopkins University, city=Baltimore, state=MD, country=USA
[…]
[DKFZ1]organization=Division of Medical Image Computing, German Cancer Research Center (DKFZ), city=Heidelberg, country=Germany
[…]
[DKFZ2]organization=Helmholtz Imaging, German Cancer Research Center (DKFZ), city=Heidelberg, country=Germany
[…]
[SMILE]organization=Smart Medical Imaging, Learning and Engineering (SMILE) Lab, Medical UltraSound Image Computing (MUSIC) Lab, School of Biomedical Engineering, Shenzhen University Medical School, Shenzhen University, city=Shenzhen, country=China
[…]
[MedIG]organization=Institute of Biomedical and Health Engineering, Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences, city=Shenzhen, country=China
[…]
[LEM]organization=Pattern Recognition Lab, Friedrich-Alexander-University Erlangen-Nuremberg, city=Erlangen, country=Germany
[…]
[AGH1]organization=Department of Measurement and Electronics, AGH University of Krakow, city=Krakow, country=Poland
[…]
[AGH2]organization=MedApp S.A., city=Krakow, country=Poland
[…]
[XMU]organization=National Institute for Data Science in Health and Medicine, Xiamen University, city=Xiamen, country=China
[…]
[Sano1]organization=Faculty of Informatics, Mathematics, and Mechanics, University of Warsaw, city=Warsaw, country=Poland
[…]
[Sano2]organization=Faculty of Electronics and Information Technology, Warsaw University of Technology, city=Warsaw, country=Poland
[…]
[Sano3]organization=Sano Center for Computational Medicine, city=Krakow, country=Poland
[…]
[UCLA]organization=Department of Radiation Oncology, University of California, Los Angeles, city=Los Angeles, state=CA, country=USA
[…]
[USTC]organization=School of Biomedical Engineering, Division of Life Sciences and Medicine, University of Science and Technology of China, city=Hefei, country=China
[…]
[JST]organization=Department of Orthopaedics and Traumatology, Beijing Jishuitan Hospital, city=Beijing, country=China
supportedsupportedfootnotetext: This work was supported by Nation Natural Science Foundation of China (NSFC6247010104), National Key Research and Development Program of China (2022YFC2504304), Natural Science Foundation of Beijing (L222136), and Beijing Science and Technology Project (Z221100003522007 and Z241100009024030)
1 Introduction
Pelvic fractures are among the most severe orthopedic injuries, often resulting from high-energy trauma such as motor vehicle accidents and falls from height. These fractures are associated with a disability rate exceeding 50% and a mortality rate above 13%, making them the deadliest of all compound fractures (Kowal et al., 2007; Sugano, 2013). The complex anatomy of the pelvis, combined with surrounding soft tissues, makes surgical intervention particularly challenging. In recent years, robotic-assisted closed fracture reduction surgeries have demonstrated improved clinical outcomes with reduced bleeding, lower infection rates, decreased vascular and nerve damage, and faster postoperative recovery (Ge et al., 2022). The precision and outcome of these image-guided robotic surgery heavily rely on accurate preoperative and intraoperative imaging and segmentation.
Automated pelvic fracture segmentation is a crucial step in trauma diagnosis and surgical planning. In 3D CT scans, precise fracture segmentation is a prerequisite for fragment-wise reduction planning, enabling accurate pose and trajectory planning, as well as screw planning, for surgical realignment and fixation (Liu et al., 2025; Shi et al., 2025). It also plays a key role in and supports detailed fracture classification, aiding in clinical decision-making. In 2D C-arm X-ray imaging, fracture segmentation is equally important for diagnostic interpretation and intra-operative guidance. By segmenting fractures in radiological images, surgical plans can be accurately transferred to the operating space, facilitating C-arm imaging-based navigation and registration (Han et al., 2021).
Traditionally, fracture segmentation in CT has been performed semi-automatically using 3D interactive tools. In the current surgical planning workflow for robotic pelvic fracture reduction, the fractured pelvis is typically extracted using a graph-cut algorithm, with manually selected foreground and background seed points. Following this step, each bone fragment must be individually isolated. For completely separated fragments, connected component analysis can be used for extraction. However, when fragments remain connected or collided, operators must carefully delineate the fracture surfaces, either by manually tracing contours slice by slice in 2D images or by examining the fracture structure in a 3D view. This process is highly labor-intensive, often requiring 0.5 to 1 hour for complex fractures with overlapping surfaces, not to mention the additional time required for subsequent fragment pose planning. The efficiency of this workflow is particularly critical, as pelvic fracture reductions are typically non-elective procedures, sometimes required in emergency settings. Moreover, the segmentation accuracy is heavily dependent on the skill and experience of the operator.
For intraoperative X-ray imaging, fracture segmentation holds significant potential to facilitate pelvic reduction surgery; however, it has not yet been adopted into clinical routine. Pelvic fracture fixation is a highly complex procedure that relies on radiographic imaging to accurately place Kirschner wires and orthopedic screws within specific bony corridors, following a detailed preoperative plan. In intraoperative cone-beam CT (CBCT) imaging, efforts have been successfully made to automatically extract bone regions and register them with preoperative CT scans (Liu et al., 2024). However, similar techniques for X-ray imaging remain underexplored. While C-arm X-ray imaging is far more widely used than CBCT, its segmentation process presents several challenges, including the difficulty in distinguishing overlapping pelvic structures, the limited visibility of fracture lines, and the need for precise alignment of fractured fragments. These challenges highlight the need for improved segmentation techniques in X-ray imaging to enhance intraoperative guidance with lower radiation exposure and wider applicability.
Automating pelvic fracture segmentation remains an unsolved challenge due to the diverse shapes and positions of bone fragments and the complex fracture surfaces caused by bone collisions. Despite the success of deep learning in various medical imaging applications, its use in automated pelvic fracture segmentation remains relatively limited. This is primarily attributed to the lack of extensive image datasets and annotations specific to fractured pelvises.
The Pelvic Bone Fragments with Injuries (PENGWIN) Challenge 2024 was launched to address these challenges by fostering the development of robust, automated segmentation techniques for pelvic fractures in both 3D CT and 2D X-ray modalities. Hosted as a satellite event at the 27th International Conference on Medical Image Computing and Computer Assisted Intervention (MICCAI 2024), this challenge aimed to:
  1. 1.
    Advance the state of the art in pelvic fracture segmentation using deep learning.
  2. 2.
    Provide a high-quality, publicly available dataset to facilitate further research in medical image analysis and orthopedics.
  3. 3.
    Promote interdisciplinary collaboration between researchers in medical imaging, artificial intelligence, and orthopedic surgery.
To achieve these goals, we curated a comprehensive, multi-institutional dataset consisting of CT scans from 150 patients scheduled for pelvic reduction surgery. These scans were acquired from multiple research centers using various scanning devices to ensure diversity in imaging conditions and patient cohorts. The ground-truth segmentations of the hipbones, sacrum, and fracture fragments were semi-automatically annotated and validated by medical experts. In addition, we generated 48,600 high-quality, realistic X-ray images using the DeepDRR method (Unberath et al., 2018), simulating diverse virtual C-arm angles and incorporating surgical tool interference to closely reflect real-world intraoperative conditions. Participants were challenged to develop fully automated segmentation algorithms for 3D CT and 2D X-ray tasks, contributing novel methods that improve the accuracy, robustness, and clinical applicability of fracture segmentation.
In this paper, we present an in-depth review of the challenge, providing a comprehensive summary of related challenges and studies, participation statistics, algorithmic approaches, comparative performance analysis, and key insights gained.
2 Related works
2.1 Orthopedic image processing challenge
Table 1: Summary of major orthopedic image processing algorithmic challenges between 2017 and 2024. Challenges directly related to bone fractures are shown in bold.
Acronym Challenge title Organization Year Modality Anatomy Task Data size
PBA-RSNA (Halabi et al., 2019) RSNA pediatric bone age challenge RSNA 2017 X-ray Hand Scoring 14,236
AASCE (Wang et al., 2021) Accurate automated spinal curvature estimation MICCAI 2019 X-ray Spine Measurement 707
VerSe’19 (Sekuboyina et al., 2021) Large scale vertebrae segmentation challenge 2019 MICCAI 2019 CT Spine Identification, segmentation 160
VerSe’20 (Sekuboyina et al., 2021) Large scale vertebrae segmentation challenge 2020 MICCAI 2020 CT Spine Segmentation 374
RA2-DREAM (Sun et al., 2022) Automated scoring of radiographic joint damage DREAM 2020 X-ray Hand, foot Scoring 698
object-CXR Automatic detection of foreign objects on chest X-rays MIDL 2020 X-ray Chest Identification 10,000
KNOAP2020 (Hirvasniemi et al., 2023) Knee osteoarthritis prediction challenge - 2020 MRI, X-ray Knee Classification 423
RibFrac (Yang et al., 2024) Rib fracture detection and classification challenge MICCAI 2020 CT Rib Identification, classification 660
AutoImplant (Li et al., 2021) Automatic cranial implant design challenge MICCAI 2020 Binary volume Cranium Implant design 200
AutoImplant2021 (Li et al., 2023) Automatic cranial implant design challenge 2021 MICCAI 2021 Surface model Cranium Implant design 681
CSF-RSNA (Hu et al., 2024) RSNA cervical spine fracture AI challenge RSNA 2022 CT Cervical spine Identification 1,400
K2S (Tolpadi et al., 2023) From undersampled k-space to automatic segmentation MICCAI 2022 MRI Knee Segmentation 350
FBD-Kaggle (Khanal et al., 2023) Fractured bone detection challenge - 2023 CT Limb Identification 24
CL-Detection (Wang et al., 2016) Cephalometric landmark detection in lateral X-ray Images MICCAI 2023 X-ray Skull Identification 600
LSDC-RSNA (Richards et al., 2024) RSNA lumbar spine degenerative classification AI challenge RSNA 2024 MRI Lumbar spine Classification 1,975
SPIDER (van der Graaf et al., 2024) Spine segmentation: discs, vertebrae, and spinal canal - 2024 MRI Lumbar spine Segmentation 544
We review the algorithmic challenges related to orthopedic image processing that took place between 2017 and 2024. Our search scope includes: (1) challenges endorsed by major medical imaging conferences, such as MICCAI, the IEEE International Symposium on Biomedical Imaging (ISBI), and Medical Imaging with Deep Learning (MIDL); and (2) publicly accessible challenges from key online platforms, including the Grand Challenge platform, Dialogue for Reverse Engineering Assessments and Methods (DREAM) Challenges, Kaggle Competitions, and Radiological Society of North America (RSNA) Challenges. This list of registries was compiled with reference to the MICCAI Special Interest Group (SIG) for Challenges. The 16 challenges surveyed are summarized in Table 1. Furthermore, challenges that are directly related to bone fractures are reviewed in detail.
The RibFrac Challenge (Yang et al., 2024) was launched with MICCAI 2020, aimed to advance automated rib fracture detection and diagnosis. The dataset comprises over 5,000 rib fractures across 660 CT scans. The challenge included two tasks. In the first task, fracture detection was formulated as an instance segmentation problem, with annotated fractured regions provided as ground truth. Due to the inherent ambiguity in delineating fracture boundaries, the segmentation outputs were primarily used to evaluate detection performance, with the free-response receiver operating characteristic (FROC) analysis employed for assessment. In the second task, the detected rib fractures were further classified into four clinically relevant categories: buckle fractures, non-displaced fractures, displaced fractures, and segmental fractures. Classification performance was evaluated using the F1 score.
The 2022 RSNA Cervical Spine Fracture AI Challenge (Hu et al., 2024) focused on rapid detection and localization of cervical spine fractures, which is critical to prevent neurological deterioration and paralysis following trauma. The dataset includes over 3,000 CT scans, with an approximately equal distribution of positive and negative cases. Participants were tasked with predicting the presence or absence of fractures at each of the seven cervical vertebrae, but were not required to locate them within the CT volume. Submissions were evaluated based on a weighted multi-label logarithmic loss, assessing the predicted probability of fracture at each vertebra.
The Fractured Bone Detection Challenge (Khanal et al., 2023) was a lightweight, independent challenge launched on Kaggle. It provided a small dataset of 24 CT scans covering both upper and lower limbs. The objective was to identify fractures in 2D CT slices to facilitate rapid diagnosis. Each CT slice was binary classified based on the presence of fractures, with area under the sensitivity-specificity curve (AUC) used as the evaluation metric.
In summary, challenges focusing on orthopedic fractures remain limited in number, with no prior challenge specifically targeting pelvic bone or X-ray segmentation. Existing efforts have predominantly focused fracture detection and localization at a coarse level. However, the emerging adoption of image-guided reduction surgery highlights the growing demand for voxel-wise fracture fragment segmentation, calling for new challenges and benchmarks to drive technical advancements in this domain.
2.2 Automated fracture segmentation
2.2.1 Fracture segmentation in CT
Traditional non-learning-based methods for automated and semi-automated CT fracture segmentation have been comprehensively summarized in previous studies (Liu et al., 2025). These approaches include intensity thresholding, graph cut, and region growing methods. While such methods can be effective in cases involving non-contact fractures, they often struggle in more complex scenarios, such as fractures with fragment contact, impaction, or compression (Krčah et al., 2011). In contrast, deep learning-based fracture segmentation remains a relatively underexplored area; however, recent studies have demonstrated its significant potential. In the following, we review the latest advancements in learning-based approaches for automated fracture segmentation in CT.
Early studies have explored the direct application of established segmentation networks to fracture fragment segmentation in CT scans. Yang et al. (2022) applied a Mask R-CNN-based instance segmentation model to locate and extract individual fragments from 2D CT slices of intertrochanteric fractures. Kim et al. (2023) utilized DeepLab v3+ for semantic segmentation of tibial and fibular fracture fragments in 2D CT slices. In their approach, fragment labels were manually assigned during data preparation; however, this strategy raises concerns regarding robustness due to potential label inconsistencies across patients and the limited scalability of treating individual fragments as semantic classes. Wang et al. (2022a) adopted a 3D V-Net model to segment intertrochanteric femoral fractures, using a three-class labeling scheme to distinguish the proximal femur, fracture fragment, and distal femur.
In our previous study, FracSegNet, fracture segmentation is performed using a two-stage pipeline comprising two sequential networks (Liu et al., 2023). First, an anatomical segmentation network extracts the hipbones and sacrum from CT images. Subsequently, a fracture segmentation network differentiates the primary fragments (label 1) and secondary fragments (label 2) within each bone region. The fracture segmentation network is trained using a distance-weighted loss, which emphasizes attention near fracture boundaries and improves performance in these critical regions. This approach enables semantic segmentation of both primary and secondary fragments, thereby avoiding the need to explicitly handle a variable number of fragments during network inference. Following segmentation, secondary fragments are further separated using connected component analysis. However, this post-processing step is insufficient when secondary fragments are in contact, leading to failures in accurately separating adjoining fragments.
Building upon the same primary-secondary representation, Zeng et al. (2024) proposed a pelvic fracture segmentation network that incorporated the prediction of a fragment prior distance map as an auxiliary task to better capture the spatial relationships between fragments. In addition, a dynamic multi-scale feature fusion module was introduced to address the challenge posed by the varying sizes of fragments.
In addition, Deng et al. (2023) explored the synergistic integration of fragment segmentation and reduction pose planning for femoral fractures. Their method takes as input segmented bone surface point clouds from both the fractured and contralateral healthy sides. By using the healthy side as a template, the segmentation of fracture fragments is refined through an alternating process that minimizes the distance between the current fracture model and the template. However, this approach is only applicable when a complete contralateral template is available and may be limited by natural anatomical asymmetry.
2.2.2 Fracture segmentation in X-ray
Early applications of deep learning for fracture analysis in X-ray imaging primarily focused on fracture detection and classification, with some approaches incorporating bone segmentation as a pre-processing step (Joshi and Singh, 2020).
Fracture detection can be performed at the image level, where a model determines the presence or absence of fractures. For instance, Cheng et al. (2019) developed a convolutional network to detect hip fractures from plain frontal pelvic radiographs, achieving an accuracy of 91%. To enhance interpretability, they further applied gradient-weighted class activation mapping (Grad-CAM) to visualize fracture regions, yielding an accuracy of 95.9% for lesion identification. Similar binary fracture classification approaches have been studied on other anatomies like long bones and spine (Wang et al., 2022b).
Beyond image-level classification, fractures can be localized more precisely using bounding boxes or segmentation contours. This is particularly valuable for hairline or incomplete fractures, where the fracture line is clearly visible in the radiograph, though the entire fragment cannot be isolated. Traditional image processing and machine learning-based methods have been explored for fracture detection using bounding box representations (Myint et al., 2018; Hržić et al., 2019; Hardalaç et al., 2022). For example, Joshi et al. (2022) employed a feature pyramid network with transfer learning to perform both fracture detection (bounding boxes) and segmentation (contours) in wrist bone fractures. Turk et al. (2024) investigated the effect of different backbone architectures on fracture segmentation in X-ray images. Their study utilized a dataset of 4,083 scans covering hand, leg, hip, and shoulder fractures, with 4,497 distinct fracture sites. The highest-performing method achieved an Intersection over Union (IoU) exceeding 0.93.
Additionally, in the pelvic region, a recent study proposed the use of deep networks to segment the entire pelvis in X-ray images in the presence of fractures, demonstrating the potential of deep learning methods in handling fracture cases (Lee et al., 2024). In spinal X-ray imaging, efforts have also been made to detect, segment, and assess vertebral compression fractures. Since spinal fractures are often non-displaced and lack clearly visible fracture lines, existing segmentation methods have primarily focused on extracting the full outline of the affected vertebrae (Kim et al., 2021).
In short, similar to the related challenges discussed in Sec. 2.1, existing studies on X-ray fracture analysis have prioritized fracture detection, with some leveraging deep learning-based segmentation as a means to refine localization. While segmentation provides a more precise localization than bounding boxes to support trauma diagnosis, current methods lack the ability to delineate fracture lines accurately or extract individual fractured fragments. Bone fragment segmentation in X-ray images remains particularly challenging due to fragment overlap—especially in the pelvis, where the ring structure complicates separation—and the presence of soft tissues in the projection image, which obscures boundaries and limits the effectiveness of standard segmentation approaches.
3 Challenge description
3.1 Challenge setup
The PENGWIN Challenge 2024 is a one-time event with fixed submission deadline comprising two independent tasks, with participants free to compete in either or both. Each task included a development phase and a final test phase, following a structured timeline. The training data was released on April 5, 2024, enabling participants to develop and refine their models. The validation phase, beginning on May 15, 2024, allowed teams to test their models on a separate validation set multiple times, with the results displayed on a public leaderboard. In the final test phase, participants submitted their models as Docker containers on the Grand Challenge platform [^1] Challenge website: https://pengwin.grand-challenge.org/ by August 19, 2024. The final results were announced on August 31, 2024, and the top three teams in each task were awarded cash prizes and certificates, conditional upon the open-sourcing of their code in a public repository. The top three teams were invited to present their methods at MICCAI 2024 in October. Only the members of the organizing team have access to the test cases. Members of the organizers’ institutes could participate but were not eligible for awards. Participating teams could publish their own results separately, with a one-year embargo period.
In the CT task, participants were provided with 3D pelvic CT scans containing fractures. The expected output was a voxel-wise segmentation mask, labeling the sacrum, hipbones, and individual bone fragments within the scan.
In the X-ray task, participants received 2D simulated X-ray images derived from CT scans. The objective was to generate a pixel-wise segmentation mask, identifying the sacrum, hipbones, and bone fragments in each X-ray image. Note that in this separate task, participants did not have access to the corresponding CT image for running the X-ray algorithms.
Participants were permitted to use publicly available datasets and pre-trained network models, provided these were under a permissive license. Templates for preparing a container submission and evaluation scripts were provided to participants before the validation phase. The online platform allowed, at the maximum, a runtime environment with a NVIDIA T4 Tensor Core GPU, eight vCPUs, 32 GB of memory, and 225 GB NVMe SSD for running the container. The time limit was set to 10 minutes per case, including the container loading, image loading, and result saving times.
3.2 Data
3.2.1 Curation of the CT dataset
The CT data were collected from 150 patients who were planning to undergo pelvic reduction surgery between 2017 and 2023 from six medical centers in China, including Beijing Jishuitan Hospital (JST), Foshan Hospital of TCM (FSHTCM), the First Bethune Hospital of Jilin University (JLUFH), the Third Bethune Hospital of Jilin University (JLUTH), Nanfang Hospital (NFH), and Tianjin Hospital (TJH). The CT scans were acquired from a diverse array of machines symmarized in Table 2. The retrospective data usage has been approved by the Ethics Committee of the institutes.
Table 2: List of CT scanners and number of cases in different centers.
Center CT Scanner Number
JST TOSHIBA Aquilion PRIME 58
UNITED IMAGING uCT 550 45
FSHTCM PHILIPS Brilliance 64 25
JLUFH SIEMENS Sensation 64 7
JLUTH TOSHIBA Aquilion ONE 7
NFH SIEMENS SOMATOM Force 1
TJH GE Optima CT660 7
Clinically used routine CT protocols have been used in each CT scanner. All scans were acquired in high quality before fracture reduction surgery and do not present significant artifacts. During the acquisition process, all patients were positioned in supine pose.
The dataset includes patients aged 16 to 94 (44.1 ± 16.8) years, with a gender distribution of 63 females and 87 males. The dataset encompasses a variety of fracture types, including pelvic ring dislocation (PRD, 5 cases), unilateral hip fracture (UHF, 54 cases), bilateral hip fractures (BHF, 31 cases), sacral fracture (SF, 5 cases), and sacral and hip fractures (SHF, 55 cases). The number of pelvic bone fragments per case ranged from 3 to 12 (5.7 ± 1.4).
The average voxel spacing is (0.83, 0.83, 0.89) mm. In cases where extra anatomies like abdomen or femoral shaft were included, the volumes were cropped to contain the pelvic region. The image dimensions average at (488, 426, 323). All the scans were de-indentified and converted to MetaImage Medical (MHA) format.
The 150 cases are split into 100 for training, 20 for validation, and 30 for testing. To ensure a balanced representation in the test set, we employed stratified random sampling based on the five primary fracture types, so that the training and test sets have the same fracture type distribution.
3.2.2 Annotation
The dataset is processed by two experienced annotators each with more than five years of medical imaging experience and a senior expert with more than 15 years of orthopedic surgical planning experience. The data annotation was structured into a four-step workflow:
  1. 1.
    Initial automatic segmentation: We employ a pre-trained segmentation network based on the nnUNet framework to produce preliminary segmentations for the hipbones and sacrum. This network was trained on roughly 1,000 pelvic CT images in the CTPelvic1K dataset, with the majority of them not presenting any fractures (Liu et al., 2021).
  2. 2.
    Manual refinement of anatomical labels: The initial anatomical labels undergo a meticulous refinement process by annotators.
  3. 3.
    Identification of fractured fragments: Leveraging the refined anatomical labels, the annotators identify and label fractured bone fragments.
  4. 4.
    Expert validation: As a final checkpoint, a senior expert rigorously reviews, modifies, and authenticates the annotated labels, ensuring their precision and consistency.
In steps 2 and 3, the annotators were instructed to annotate three anatomical labels—left hipbone, right hipbone, and sacrum—as well as the bone fragments within each bone using the 3D Slicer software. If a bone is intact and does not have any fractures, the only fragment corresponds to its anatomical mask. Fragments with a volume smaller than 500 mm3 were omitted. The ground-truth labels were also saved in MHA format.
Figure 1: Example pelvic CT scan and the annotated segmentation rendered in 3D view. The annotation follows a two-level structure: the first level indicates the anatomical origin, and the second level distinguishes individual fragment instances.
Fig. 1 illustrates an example CT scan along with its annotated segmentation mask. In this instance segmentation task, the number of fragment instances varies across cases. The label indexing follows a two-level hierarchical structure: first, fragments are classified based on their anatomical origin, and second, they are ordered by fragment size. The CT training set is publicly available in an open repository under a CC BY-NC-SA license [^2] PENGWIN CT dataset: https://doi.org/10.5281/zenodo.10927452.
3.2.3 Generation of the X-ray dataset
Figure 2: Example simulated X-ray image and the annotated segmentation.
The PENGWIN X-ray dataset comprises digitally reconstructed radiographs (DRRs) that simulate intraoperative C-arm X-ray images, including pelvic anatomy, fracture fragments, and surgical tools. To ensure robustness against surgical hardware interference, instrument poses were randomly varied across images. The X-ray images were synthesized from CT data using the DeepDRR framework, which simulates realistic X-ray projections through a combination of analytic ray-tracing and learning-based scatter estimation and noise injection techniques (Unberath et al., 2018). The generation pipeline involves the following steps:
  1. 1.
    Material decomposition is applied to the CT volume to distinguish air, soft tissue, and bone.
  2. 2.
    Forward projection is conducted using material- and spectrum-aware ray-tracing, simulating X-ray attenuation along each ray path.
  3. 3.
    Scatter estimation is performed using a deep learning-based convolutional network trained on Monte Carlo simulation data to predict scatter contributions across the projection image.
  4. 4.
    Noise injection is applied, incorporating quantum noise and electronic readout noise to approximate realistic X-ray image formation.
The virtual C-arm parameters were randomized across the training set. Each view was sampled from a uniform distribution over a solid angle extending up to 60 degrees from the vertical axis, assuming a supine patient position. The image resolution was set to (448, 448) pixels. Ground-truth 3D segmentation masks were projected using the same camera geometry to generate corresponding 2D multi-label segmentation masks. Due to the projection nature of X-rays, overlapping bone fragments result in multi-label regions, making this a multi-label instance segmentation task. An example simulated X-ray and its ground-truth label are shown in Fig. 2.
The simulated X-ray dataset was generated from the 150 CT scans described above, following the same data partitioning scheme. For the training and validation sets, 400 X-ray images were simulated per CT scan, while for the test set, 20 X-ray images were generated per CT scan, resulting in a total of 600 test cases. The images were provided in binary Tag Image File Format (TIFF), with multi-label segmentation masks encoded as single-channel binary TIFF images to comply with the challenge platform requirements. Additionally, the encoding and decoding scripts for handling segmentation files were made available to participants to ensure consistency in data processing. The X-ray training set is publicly available in an open repository under a CC BY-NC-SA license [^3] PENGWIN X-ray dataset: https://doi.org/10.5281/zenodo.10913195.
3.3 Assessment method
3.3.1 Evaluation metrics
The performance of submitted algorithms was evaluated using a set of quantitative metrics designed to assess both anatomical segmentation accuracy and fracture fragment delineation. The evaluation incorporated both overlap-based and distance-based measures to provide a comprehensive assessment.
For both the CT task and the X-ray task, the primary evaluation metrics included intersection over union (IoU), the 95th percentile Hausdorff distance (HD95), and average symmetric surface distance (ASSD). IoU was computed on solid masks, while HD95 and ASSD were computed on surface contours. The adoption of these metrics aligns with the recommendations from the metrics reloader website and has been validated in the study by Faghani et al. (2022).
IoU measures the overlap between the predicted segmentation mask $S_{p}$ and the ground truth mask $S_{g}$, and is defined as:
$\text{IoU}=\frac{|S_{p}\cap S_{g}|}{|S_{p}\cup S_{g}|},$ (1)
where $\cap$ denotes the intersection and $\cup$ denotes the union of the two masks.
HD95 quantifies the worst-case segmentation error by measuring the 95th percentile of the distances between the predicted and ground-truth segmentation contours. It is defined as:
$\text{HD}_{95}=\max\left\{\sup_{x\in C_{p}}\inf_{y\in C_{g}}d(x,y),\sup_{y\in C_{g}}\inf_{x\in C_{p}}d(x,y)\right\}_{95\%},$ (2)
where $C_{p}$ and $C_{g}$ denote the surfaces (contours) of the predicted and ground-truth segmentations, respectively, and $d(x,y)$ is the Euclidean distance between points $x$ and $y$.
ASSD computes the average bidirectional distance between the predicted and ground truth segmentation contours, ensuring a smooth and accurate segmentation. It is defined as:
$\text{ASSD}=\frac{1}{|C_{p}|+|C_{g}|}\left(\sum_{x\in C_{p}}\inf_{y\in C_{g}}d(x,y)+\sum_{y\in C_{g}}\inf_{x\in C_{p}}d(x,y)\right),$ (3)
where distances are computed symmetrically in both directions.
Evaluation was performed on both fracture segmentation and anatomical segmentation results. Fracture segmentation performance was assessed by comparing each matched fragment instance with its corresponding ground truth. Anatomical segmentation performance was evaluated by merging the predicted labels for all bone structures and comparing them to the ground-truth anatomical labels of the left hipbone, right hipbone, and sacrum. In the case of fragment instance segmentation, each ground-truth fragment was matched to the predicted fragment with the highest IoU, ensuring a one-to-one correspondence in performance assessment. In cases where a fragment was entirely missing in the prediction, HD95 and ASSD were set to the diameter and radius, respectively, of the ground-truth fragment’s bounding sphere. As a result, six evaluation metrics were used: IoU-F, HD95-F, and ASSD-F for fracture segmentation and IoU-A, HD95-A, and ASSD-A for anatomical segmentation. This metric scheme was applied to both tasks and has been made publicly available [^4] Evaluation code: https://github.com/YzzLiu/PENGWIN-example-algorithm. .
3.3.2 Ranking scheme
The ranking procedure consisted of the following steps:
  1. 1.
    Metric calculation: For each individual test case, evaluation metrics were computed for both fracture segmentation and anatomical segmentation.
  2. 2.
    Independent ranking: For each submission, the average performance across all test cases was calculated for each metric. Submissions were then ranked independently based on their performance on each metric.
  3. 3.
    Final ranking: The final rank of each submission was determined by computing the average of its ranks across all evaluation metrics. This approach ensured that submissions were assessed holistically rather than being overly influenced by a single metric.
  4. 4.
    Tie-breaking: In the event of a tie in the final ranking, the container execution time of the submission was used as a tie-breaker, with faster models receiving a higher rank.
4 Results
4.1 Participation
The challenge attracted interest from the medical imaging and machine learning communities. A total of 239 participants from 25 countries registered for the challenge, gaining access to the dataset. Among them, 26 registered teams were formed.
During the preliminary development phase, participants actively engaged with the challenge, resulting in 256 submissions for the CT task and 84 submissions for the X-ray task.
During the final test phase, 11 teams from five countries submitted models for the CT task, while five teams from three countries participated in the X-ray task. Among them, two teams chose to complete both tasks.
4.2 Algorithm summary
In this section, we summarize the methods employed by the top five performing teams in the CT task and the top four performing teams in the X-ray task, as subsequent teams exhibited significantly larger performance gaps in comparison.
4.2.1 CT segmentation
Table 3: Summary of the top-performing participating algorithms for CT segmentation. Spatial augmentation includes random scaling, mirroring, and flipping. Intensity augmentation includes noising, blurring, and histogram adjustments.
Team External resource Anatomical network Fracture network Pre-processing Augmentation Interpolation Loss
MIC-DKFZ - UNet with residual UNet with residual standard nnUNet spatial, intensity, low-res cubic; anisotropic Dice, CE
SMILE - UNet UNet standard nnUNet spatial, cropping linear; isotropic Dice, CE
MedIG TotalSegmentator Pre-trained UNet NexToU standard nnUNet spatial, intensity, low-res cubic; anisotropic Dice, CE
MedApp-AGH FracSegNet UNet UNet standard nnUNet spatial cubic; isotropic Dice, CE
Sano - UNet UNet normalize, reorient spatial, intensity linear; isotropic Dice, CE
MIC-DKFZ (1st place, Johannsen et al.)
The authors formulated the task as two semantic segmentation problems. The first network performed anatomical segmentation of the three bone regions. The second network predicted the adaptive border boundary core (ABBC) representation, which consists of four semantic classes: background, boundary, border, and core. The core and boundary classes together define individual fragments, with the boundary class acting as a separator between the core and the background or adjacent fragments. The authors introduced a novel dynamic boundary thickness adjustment using the medial axis transform, ensuring that thin fragments remain connected within the core region. The border class was defined as the intersection of the dilated outlines of neighboring fragments. Given a prediction, connected component analysis was applied to the core class to extract initial fragment instances. The boundary class was then merged with the nearest seed to form fragment instances. To ensure anatomical consistency, fragment instances were adjusted by splitting them along anatomical borders and merging smaller fragments with the closest anatomically correct instance. Additionally, the border class was used to refine the segmentation by modifying fragments through split-merge operations, ensuring alignment with the network predictions. [^5] MIC-DKFZ CT algorithm: https://codebase.helmholtz.cloud/hi-dkfz/applied-computer-vision-lab/challenges/abbc
SMILE (2nd place, Yue et al.)
The authors proposed a two-stage segmentation framework. The first stage focused on localizing the three bone regions. To achieve this, the authors trained a U-Net model on downsampled CT scans, which reduced the computational costs and provided a global context that prevented excessive false positives when processing smaller patches. The localized regions were then refined at full resolution for accurate anatomical segmentation. The second stage focused on fracture segmentation within each of the three bone regions. The authors categorized bone fragments into primary (largest fragment) and secondary (smaller fragments) components. After obtaining initial predictions, connected component analysis was applied to reassign secondary fragment labels in accordance with evaluation rules. The final segmentation result was generated by combining the outputs from both stages. A final refinement step ensured that the fracture segmentation was aligned with the foreground mask from the first stage. [^6] SMILE CT algorithm: https://github.com/yuepeiyan/PENGWIN_Challenge
MedIG (3rd place, Pan et al.)
The authors employed a two-stage segmentation pipeline to accurately localize and segment fractured bone fragments. In the first stage, the TotalSegmentator (Wasserthal et al., 2023) model was used to localize the pelvic region, after which the three bone regions were cropped from the original CT scans. In the second stage, the NexToU model was applied to segment primary and secondary fragments within each cropped bone region. The largest connected component was extracted as the primary fragment, and the remaining secondary fragments were expected to be naturally separated. A connected component analysis step was used to isolate individual secondary fragments. [^7] MedIG CT algorithm: https://github.com/pzhhhhh2263/MICCAI-Challenge-PENGWIN2024
MedApp-AGH (4th place, Jurgas et al.)
The authors developed a refinement-based segmentation pipeline, integrating anatomical segmentation with fracture surface segmentation to ensure anatomically consistent fracture identification. The approach utilized two deep networks. The first network, adapted from the FracSegNet study (Liu et al., 2023), was responsible for segmenting the three bone regions. The second network, trained on fracture surfaces extracted from the PENGWIN dataset, identified fracture boundaries. For training, the authors generated a complete anatomical segmentation of the pelvis by merging all bone fragments. They then extracted contours of both individual fragments and the anatomical segmentation, removing overlapping edges to isolate fracture surfaces—where fractured bones were in contact. The fracture network was trained to predict these fracture surfaces from CT scans. In post-processing, segmentations from both networks were merged. A continuity analysis was then performed to separate individual fractures, with fragments labeled according to their anatomical overlap. [^8] MedApp-AGH CT algorithm: https://github.com/Jarartur/pengwin-challenge-submission
Sano (5th place, Płotka et al.)
The authors proposed PeFreCT, which employed two UNet models. The first network performed anatomical segmentation, while the second network focused on fracture segmentation, detecting fracture regions based on textural features. For training the anatomical network, the authors cropped the CT scans by first detecting the body using radio-density thresholding and removing artifacts with morphological operations. The dataset was doubled by flipping images along the sagittal plane and mirroring hipbone labels. For training the fracture network, they precomputed the fracture region mask by dilating the masks of all fragments and identifying overlapping borders to form a final binary fracture label. Since fracture detection based on texture is predominantly local, a (64, 64, 64) patching scheme was applied for both training and inference to enhance efficiency. In post-processing, the binary fracture mask were subtracted from anatomical masks, allowing for fragment separation with connected component analysis. Finally, the fragment masks were expanded to include the fracture regions, using the watershed algorithm guided by raw fracture probability maps to precisely delineate the borders between colliding fragments.
Summary
All five teams employed a two-stage approach, first performing anatomical segmentation before segmenting fracture fragments. The key variations among their methods lay in the fracture segmentation step—whether to extract individual fragments directly or to first detect border regions and infer fragment separation. Most teams adopted the nnU-Net framework, owing to its robustness and user-friendly data processing pipeline. Additionally, Dice and cross-entropy (CE) loss were commonly used, as these loss functions have demonstrated effectiveness in various medical imaging tasks.
In Table 3, we summarize some of the main components of these algorithms, including network architecture, pre-processing, and training schemes. A detailed discussion of methodology is included in Sec. 5.1.
4.2.2 X-ray segmentation
Table 4: Summary of the top-performing participating algorithms for X-ray segmentation. Spatial augmentation includes random scaling, mirroring, and flipping. Intensity augmentation includes noising, blurring, and histogram adjustments.
Team Overall scheme External resource Network Pre-processing Augmentation Loss
SMILE Bone extraction followed by primary-secondary segmentation - UNet standard nnUNet spatial, crop Dice, CE
MedIG Direct semantic segmentation for the three bones in parallel ImageNet Unet++ with RegNetY-160 crop, normalize - Dice, CE, Focal Tversky
LME Bone extraction followed by instance segmentation COCO Swin UNETR; Mask R-CNN pad intensity Dice
luckyjing Direct instance segmentation with anatomical classification - Mask R-CNN with ResNet50 normalize flip, crop, pad CE, smooth L1, IoU
SMILE (1st place, Yue et al.)
The authors addressed the limitations of semantic segmentation in handling overlapping anatomical structures in X-ray images by developing a two-stage hierarchical segmentation strategy similar to their CT solution. In the first stage, three dedicated U-Net models were trained to segment the three bone regions separately. The isolated bone regions allowed for focused substructure analysis within each area. In the second stage, a hierarchical fragment classification approach was implemented, adapted from the authors’ CT segmentation method, with modifications specific to X-ray imaging. Separate models were employed to distinguish primary and secondary fracture fragments, mitigating the overlapping structures in X-rays. While this approach effectively reduced primary-secondary overlap, challenges remained in cases where multiple secondary fragments overlapped. For post-processing, connected component analysis was applied to relabel secondary fragments, while primary fragment predictions were refined through consistency checks with the localization masks from the first stage, ensuring that only validated foreground pixels were retained. [^9] SMILE X-ray algorithm: https://github.com/yuepeiyan/PENGWIN_Challenge
MedIG (2nd place, Pan et al.)
The authors employed a multi-model approach to process each of the three bone regions in X-ray images. Each model processed a 2D X-ray image and generated a multi-channel 3D tensor as output. The first channel represents the primary fragment, while the remaining nine channels corresponds to secondary fragments. The segmentation results from the three models were then combined to construct the final output. To optimize model performance, the authors applied a loss function composed of an unweighted combination of Dice loss, binary CE loss, and focal Tversky loss. A pre-trained ImageNet model was used to initialize the network, enhancing feature extraction capabilities. The network architecture was based on a UNet++ backbone with a RegNetY-160 encoder, which provided a balance between spatial resolution and computational efficiency. [^10] MedIG X-ray algorithm: https://github.com/pzhhhhh2263/MICCAI-Challenge-PENGWIN2024
LME (3rd place, Liu et al.)
The authors proposed a category and fragment segmentation (CFS) framework. First, three separate category segmentation networks, each based on the Swin-UNETR architecture, were used to segment the three bone regions from X-ray images. After obtaining the anatomical labels, a fragment segmentation network was applied to isolate the bone fragments within each region. The fragment segmentation step employed three separate Mask R-CNN-based networks, which performed both segmentation and object detection through bounding box prediction and classification. These networks took both the X-ray images and the binary anatomical masks as input, allowing the model to focus on specific bone regions while reducing interference from surrounding structures. During inference, the network predicted multiple bounding boxes, each associated with a confidence score. Only bounding boxes with a confidence score above 0.8 were retained for further analysis. In post-processing, the fragment segmentation masks were multiplied by the anatomical masks from the first step to ensure spatial consistency. [^11] LME X-ray algorithm: https://github.com/DaE-plz/CFSSegNet
luckyjing (4th place, Ma et al.)
The authors employed Mask R-CNN with a ResNet50 backbone network for instance segmentation. The model consists of two main stages: a region proposal network that generates candidate object regions, and a detection network that refines these regions while performing anatomical classification, bounding-box regression, and mask prediction. Mask R-CNN was employed to predict pixel-wise binary masks for each region of interest (RoI). To maintain spatial alignment between RoIs and extracted features, the network replaced RoI Pooling with RoI Align, which eliminates quantization errors by using bilinear interpolation instead of discrete binning. This significantly improved mask prediction accuracy, particularly for small objects. To address the specific challenges of pelvic fragment segmentation, the loss weights between classification and segmentation were carefully adjusted during training. Additionally, to improve segmentation accuracy, a structured loss function combining IoU and cross-entropy loss was incorporated, enhancing the model’s robustness and segmentation quality in occluded scenarios.
Summary
The solutions in the X-ray task exhibited greater diversity in their overall approach compared to the CT task. The main components of these algorithms are summarized in Table 4 and further discussed in Sec. 5.1.
4.3 Performance analysis
4.3.1 CT segmentation performance
Figure 3: Example CT segmentation results from the top-performing teams. Four representative cases are shown in (a)-(d), illustrating varying levels of complexity from easy to difficult.
Example CT segmentation results are shown in Fig. 3. The performance metrics of each algorithm submitted to the final test are summarized in Table 5 and Fig. 5.
Table 5: Quantitative results for the final test phase of the CT task. HD95 and ASSD values are reported in millimeters. The best scores for each metric are shown in bold.
Team Rank ↓ IoU-F ↑ HD95-F ↓ ASSD-F ↓ IoU-A ↑ HD95-A ↓ ASSD-A ↓
MIC-DKFZ 1 0.9296 5.866 1.843 0.9810 2.368 1.290
SMILE 2 0.9096 5.562 1.663 0.9802 2.410 1.321
MedIG 3 0.9084 6.207 1.680 0.9798 2.452 1.298
MedApp-AGH 4 0.9049 6.899 2.007 0.9783 2.398 1.385
Sano 5 0.8802 6.866 1.697 0.9491 2.522 1.343
Lee-SKKU-LG 6 0.8079 17.449 3.612 0.9718 5.283 1.576
CLone 7 0.6339 46.486 8.899 0.9332 16.119 2.730
Rzjs 8 0.6787 48.742 7.963 0.9287 18.322 2.806
Zeng-SJTU 9 0.5423 46.457 17.757 0.6235 37.966 12.306
Zou-SZU 10 0.5944 86.892 18.090 0.6708 115.814 22.005
watsons 11 0.4940 83.746 19.186 0.6165 90.970 16.478
Baseline - 0.9003 6.967 2.178 0.9804 2.423 1.299
Inter-observer - 0.9848 1.317 0.306 0.9996 1.069 0.474
Ensemble - 0.9364 4.055 1.347 0.9819 2.396 1.403
Figure 4: Distribution of segmentation errors for the top five performing algorithms on the CT test set. The first row shows the fragment-wise results. The second rows shows the anatomical results.
In addition, we evaluate the performance of our previously proposed CT semantic segmentation method, FracSegNet, described in Sec. 2.1, as a retrospective baseline. Note that the results reported in the original study were based on a different test set, and the evaluation ground truth in that work did not separate secondary fragments as required in this challenge.
4.3.2 X-ray segmentation performance
Figure 5: Example X-ray segmentation results from the top-performing teams. Four representative cases are shown in (a)-(d), illustrating varying levels of complexity from easy to difficult.
Example X-ray segmentation results are shown in Fig. 5. The performance metrics of each algorithm submitted to the final test are summarized in Table 6 and Fig. 6.
Table 6: Quantitative results for the final test phase of the X-ray task. HD95 and ASSD values are reported in pixel unit. The best scores for each metric are shown in bold.
Team Rank ↓ IoU-F ↑ HD95-F ↓ ASSD-F ↓ IoU-A ↑ HD95-A ↓ ASSD-A ↓
SMILE 1 0.7736 37.366 8.545 0.9238 13.299 2.222
MedIG 2 0.7640 40.931 9.043 0.9268 15.979 2.462
LME 3 0.7150 42.986 10.355 0.8760 21.474 4.049
luckyjing 4 0.7145 45.752 10.398 0.8125 46.060 8.877
Puvvula-NIT 5 0.5506 158.171 28.092 0.7599 150.135 23.778
Ensemble - 0.8022 30.092 6.378 0.7970 69.878 12.268
Figure 6: Distribution of segmentation errors for algorithms on the X-ray test set. The first row shows the fragment-wise results. The second rows shows the anatomical results.
Figure 7: Distribution of the IoU-F score of the five top performing teams in both tasks across cases. In the X-ray task, the scores for the 20 synthetics samples per original CT scan are averaged.
Fig. 7 presents the distribution of IoU-F across test cases for both tasks. Among them, case 19 (d in Fig. 3 and Fig. 5) emerged as an outlier, exhibiting the lowest segmentation accuracy in both tasks. This was mainly attributed to the presence of multiple small fragments, particularly in the ramus pubis region.
To assess the statistical significance of pairwise differences between competing teams, we employed the Wilcoxon signed-rank test on the IoU-F scores. We specifically conducted a one-sided Wilcoxon signed-rank test to determine whether the IoU-F scores of one team were statistically significantly higher than those of another. In this setting, the null hypothesis states that the median difference between paired observations is zero or negative. The alternative hypothesis asserts that the IoU-F scores of the first team are systematically greater than those of the second team. The resulted significance matrices are shown in Fig. 8. A p-value below 0.05 was considered statistically significant.
Figure 8: Significance matrices showing the p-values of the participating algorithms using one-sided Wilcoxon signed-rank tests.
4.4 Inter-observer agreement
In order to assess the quality of the ground truth, as well as to compare the algorithm’s performance against human experts’, we perform a retrospective analysis to assess inter-observer variability by manually annotating the CT test set with an additional independent annotator, using the same protocol described in Sec. 3.2.2.
The assessment results are also summarized in Table 5. The high agreement between annotators confirmed the reliability of the ground truth. The primary source of variation stemmed from differences in determining whether an incomplete fracture should be labeled as two separate fragments or a single fragment in certain cases.
For the X-ray task, however, a direct evaluation of inter-observer variability is not feasible, as the ground-truth labels is automatically generated through projection, and manual annotation of the projected labels proved exceedingly difficult.
4.5 Impact of fragment size and count on segmentation
We investigated the impact of fragment size and fragment count on segmentation performance, using results from the top five performing algorithms in both the CT and X-ray tasks. The findings are illustrated in Fig. 9.
Overall, IoU-F scores decrease as fragment size decreases and fragment count increases, indicating the increased difficulty in accurately isolating or delineating small fragments, particularly in complex fracture scenarios.
Figure 9: Impact of fragment size and count on the IoU-F metric for the top five performing algorithms in both tasks. In (a) and (c), fragment sizes are binned in intervals of 500,000 mm3 for CT and 10,000 pixel3 for X-ray to enhance visual clarity.
4.6 False-positive analysis
Table 7: False-positive predictions in the final test phase. The mean and standard deviation of unmatched fragment counts are reported. The upper and lower sections show results for the CT and X-ray tasks, respectively.
Team Rank Mean Std
MIC-DKFZ 1 0.53 0.78
SMILE 2 0.20 0.41
MedIG 3 0.93 0.94
MedApp-AGH 4 0.50 0.68
Sano 5 19.23 3.74
SMILE 1 0.50 0.95
MedIG 2 0.42 0.67
LME 3 3.12 1.13
luckyjing 4 2.11 1.64
In our evaluation framework, false-positive (FP) fragments were not explicitly penalized at the instance level, but were instead indirectly reflected through the anatomical segmentation metrics. To better understand the prevalence and cause of FPs, we additionally computed the number of unmatched predicted fragments for each participating algorithm with respect to the ground truth, regardless of its size or shape. This analysis provides insight into over-segmentation tendencies and complements the primary evaluation metrics. The results are summarized in Table 7.
In the CT task, team SMILE achieved the lowest FP rate, attributable to its coarse-to-fine anatomical segmentation strategy, which effectively excluded non-pelvic bony structures, and the use of a primary-secondary representation scheme that tended to under-segment small fragments rather than over-segment them. The algorithm developed by team Sano exhibited a higher FP rate due to its tendency to over-segment small fragments. However, a portion of these fragments actually appeared in plausible locations—such as regions with thin, deformed cortical bone—and could be eliminated through post-processing with size-based filtering when needed. Other sources of FP predictions included erroneous anatomical segmentation that mistakenly included patients’ hands or the CT table, as well as over-detection in regions where fractures commonly occur.
In the X-ray task, the primary-secondary semantic segmentation approach adopted by team SMILE and team MedIG resulted in fewer FPs compared to the instance segmentation strategies employed by team LME and team luckyjing. While instance-based methods hold potential for more precise identification of small fragments, they tend to be less robust in practice.
4.7 Ranking robustness
Figure 10: Bootstrap-based stability analysis of team rankings in both tasks. Each dot represents the frequency with which a team achieved a specific rank across 1,000 bootstrap samples of the test set. Dot size is proportional to the frequency of that rank being assigned.
To evaluate the robustness of the final team rankings against variations in the test set, we performed a bootstrap analysis using 1,000 randomly resampled replacements (Wiesenfarth et al., 2021). For each bootstrap sample, the final ranks were recalculated following the same evaluation protocol. We then calculated the Kendall rank correlation coefficient $\tau$ between each bootstrap-derived ranking and the original ranking computed on the full test set, and visualized the ranking variability of each team using a rank stability plot.
In the CT task, the bootstrap analysis yielded a mean Kendall’s $\tau$ of 0.9163, with a 95% confidence interval of [0.807, 1.000], indicating a good agreement between the bootstrap-derived rankings and the original ranking. The X-ray task exhibited even higher stability, with a mean Kendall’s $\tau$ of 0.9997 and an extremely narrow 95% confidence interval of [0.999, 1.000]. Fig. 10 illustrates the robustness and variability of the ranking outcomes, with larger and more concentrated dots indicating greater ranking stability.
4.8 Ensemble of participating algorithms
We investigated the feasibility of ensembling multiple segmentation algorithms to achieve improved performance and robustness.
In the CT task, to obtain a consensus segmentation from multiple team submissions, we developed a graph-based ensemble strategy tailored for instance-level fracture segmentation. This approach was applied independently to each of the three anatomical regions and incorporates spatial overlap analysis, voting-based fusion, and standardized label assignment.
Specifically, for each test case, we first extracted all candidate fragment masks from the submissions of the top five teams. Candidate fragments were grouped by anatomical region using predefined label ranges. To identify similar fragments across different submissions, we constructed a pairwise similarity graph where each node corresponds to a fragment and edges connect pairs with an IoU greater than 0.5. Clusters were then identified as the connected components of this graph, each representing a consensus fragment candidate. Within each cluster, a majority voting scheme was applied to the set of masks to generate a single fused segmentation. Clusters with fragment size less then 500 mm3 were excluded. To address overlapping masks, fragments were written in order of size, allowing smaller fragments to overwrite larger ones and thus better preserve detail in thin regions.
As shown in the last row of Table 5, the method achieved superior fracture segmentation performance across all three metrics compared to the participating algorithms, while attaining anatomical segmentation performance comparable to that of the top-performing method.
In the X-ray task, the projection nature of the images introduces ambiguity, making it difficult to determine whether two overlapping fragments from different algorithms represent the same structure or distinct fragments that overlap in the projection. To mitigate this, we adopted a conservative merging strategy of the four teams: all candidate fragments were initially preserved, and only those with a high degree of spatial agreement—defined by an IoU greater than 0.9—were merged. While this approach helped avoid erroneous merging of distinct structures, it also increased the risk of including FPs or redundant, highly similar fragments in the final output.
As shown in the last row of Table 6, the ensemble method leveraged the diversity of participating algorithms to achieve improved fragment-wise segmentation performance. However, this gain came at the cost of a notable drop in anatomical segmentation accuracy, due to an increased number of FP predictions.
5 Discussion
5.1 The overall scheme for fragment segmentation
In the CT task, two-stage approaches consisting of anatomical segmentation followed by fragment segmentation were widely adopted by the participants. This two-stage strategy was chosen for multiple reasons: (1) The sacroiliac joint, whether dislocated due to trauma or not, exhibits distinct characteristics compared to fracture surfaces within individual pelvic bones, making it difficult to effectively address both in a single-step segmentation approach; (2) The effectiveness of this scheme has been previously validated in the FracSegNet study, where bone region extraction simplified the task; (3) Pre-trained large models, such as TotalSegmentator and the CTPelvic1K baseline model, could be directly utilized to extract the bone regions; and (4) The evaluation metric design explicitly required accurate anatomical segmentation as a distinct criterion, ensuring alignment with the requirements of subsequent surgical planning procedures (Liu et al., 2025). We computed the correlation between IoU-A and IoU-F in the CT task across all participating teams. A strong linear relationship was observed, with Pearson’s correlation coefficient of 0.870 (p = 0.0005), suggesting that accurate anatomical segmentation serves as a critical foundation for successful fracture segmentation.
The primary distinction among the top teams lay in their fracture segmentation strategies in the second stage. To convert the instance segmentation problem, which involves a variable number of fragments, into a more tractable semantic segmentation problem, some teams adopted a primary-secondary representation, while others employed a boundary-core representation to isolate individual fragment instances.
The primary-secondary representation can be challenging when the definition of the ”primary” fragment is ambiguous or when multiple secondary fragments collide in adjacent regions, a scenario frequently observed in anterior pelvic ring fractures. In contrast, the boundary-core representation is well-suited for handling such cases but depends heavily on precise boundary segmentation. Small errors in boundary prediction can lead to significant connectivity changes, which is a potential concern when dealing with incomplete fractures.
The solutions in the X-ray task exhibited greater methodological diversity compared to the CT task. While some teams employed two-stage approaches similar to those used in the CT task, others opted for more direct methods, leveraging advances in 2D image detection and segmentation techniques, such as Mask R-CNN. These approaches, which have demonstrated strong performance across various applications, were adapted to accommodate multi-class overlapping labels, addressing the unique challenges of the X-ray segmentation task.
All of the four methods summarized in this task focused on fragment labels rather than fracture lines, as overlapping fragments remain inseparable due to the projection nature of X-ray imaging, even when the fracture regions are removed. The primary-secondary representation, despite being adopted by both the first- and second-place teams, exhibited a notable limitation: small fragments classified as secondary were more frequently connected compared to the CT images, making them difficult to separate through post-processing. This challenge posed a significant drawback in accurately delineating individual small fragments with semantic segmentation approaches.
The results of X-ray segmentation algorithms proved insufficient to reliably support clinical decision-making or subsequent tasks. Several factors contributed to this limitation: (1) A single bone fragment may appear as multiple disconnected parts in the X-ray due to the limited field of view; and (2) The overlap of fragments with other tissues significantly compromises the delineation of fracture lines and hinders fragment separation.
To overcome these challenges, incorporating prior knowledge of the fragments is essential. One potential approach is to utilize multiple X-ray images from different views or targeted regions, as overlapping fragments in one projection may appear as distinct regions in a near-orthogonal plane, thereby improving segmentation accuracy with the mutually complementary information. Another promising direction involves integrating preoperative CT segmentation into the process. Given the availability of a patient’s CT scan, 2D/3D registration can be employed to estimate the approximate positions of fragments (Grupp et al., 2019). However, additional efforts are required to account for the exact C-arm geometry, which is typically unknown beforehand, and to address fragment position changes that occur between CT and X-ray acquisition, both before and during surgery. Although these methodologies extend beyond the current scope of the PENGWIN challenge, they hold significant potential for translating fracture segmentation techniques into clinical practice.
5.2 Limitations and future directions of PENGWIN
5.2.1 Dataset
Compared to related challenges organized at major medical imaging conferences, where the number of CT samples typically ranges from 160 to 1,400, the PENGWIN dataset is relatively small in size. This limitation is primarily due to the rarity of pelvic fracture CT data, as pelvic fractures account for only 3% of all fractures (Hu et al., 2023).
Additionally, while the X-ray dataset closely resembles real clinical data due to the DeepDRR simulation approach, it remains synthetic rather than derived from real patient imaging. Although DeepDRR facilitates high-fidelity X-ray generation, enabling this prospective study with a large annotated dataset, the absence of real C-arm X-ray images raises concerns regarding its direct clinical translation.
For future iterations of the PENGWIN challenge, we plan to substantially expand the fractured CT dataset, increasing the number of cases to 300–500 scans. Moreover, we aim to incorporate real clinical C-arm X-ray images in test set and annotate them using 2D/3D registration approaches, aligning them with corresponding CT scans to ensure ground-truth accuracy and facilitate clinically relevant segmentation tasks.
5.2.2 Uncertainty in annotation
As discussed in Sec. 4.4, the annotation process involved inherent uncertainty, requiring expert annotators to determine whether an incomplete fracture (partially fractured but non-displaced) should be classified as two separate fragments or a single piece. This decision was primarily based on their understanding of the subsequent surgical process, particularly whether the two segments could be manipulated as a single unit during reduction without undergoing significant deformation.
In clinical practice, this determination is made comprehensively by the surgeon, considering not only medical imaging but also physical examinations and additional clinical information, such as the mechanism of injury. Given this complexity, a collaborative interaction between human experts and segmentation algorithms may be necessary. The current approaches, particularly the primary-secondary labeling scheme, still require manual refinement by operators, which may involve merging or splitting labels. This process remains time-consuming, especially when a larger number of fracture surfaces require delineation.
For future iterations of PENGWIN, we aim to explore interactive fragment segmentation. One potential direction is to incorporate a fast manual seed point selection step, allowing operators to define the number and approximate positions of fragments before segmentation. This approach is expected to eliminate uncertainty regarding fragment count while enhancing segmentation performance by incorporating prior anatomical knowledge into the process.
Another potential direction is to integrate uncertainty quantification, allowing the network to generate multiple segmentation results that reflect different interpretations of incomplete fractures or unclear fracture lines. This approach would enable operators to select the most appropriate segmentation scheme, with their attention guided toward highlighted regions of uncertainty for further assessment.
5.2.3 Challenge logistics
Finally, we summarize several logistical limitations identified during the challenge and the post-challenge participant survey, which should be addressed in future iterations and serve as lessons for other challenge organizers to consider.
First, although the dataset and preliminary validation phase were released well in advance, the final test phase deadline was extended by three days due to multiple participant requests. To maintain fairness, teams that had already submitted their algorithms were given an opportunity to update their submissions. However, such extensions should be avoided in future iterations to uphold the rigor and integrity of the challenge timeline.
Second, apart from providing an algorithm template, we did not offer a baseline method for participants to build upon, which may have contributed to delays in submissions. Including a reference baseline in future iterations could help standardize development efforts and provide a clearer starting point for participants.
Lastly, we encountered challenges in retrospectively collecting algorithm descriptions from participants after the challenge had ended, particularly from lower-performing teams. To address this, future iterations will require participants to submit a technical description of their method along with their final submission, ensuring that essential technical details are documented in a timely manner.
6 Conclusion
In this paper, we have presented the PENGWIN challenge on pelvic fracture segmentation in CT and X-ray images, organized as part of MICCAI 2024. We provided a comprehensive review of related challenges and studies, a detailed overview of the challenge setup, dataset, participant contributions, and performance analysis.
In the CT task, the winning team achieved a high level of accuracy, with an average fragment-wise IoU of 0.930. In the X-ray task, the best-performing algorithm attained an IoU of 0.774, which, while promising, remains insufficient for reliably supporting clinical applications.
Beyond quantitative performance, the participating algorithms demonstrated great methodological diversity, offering valuable insights for future research in this emerging field. The various approaches and strategies employed have been thoroughly discussed in this paper. While the focus of this challenge was on pelvic fractures, we believe that many of the developed methodologies and key findings will generalize to other types of fractures. The pelvis presents one of the most complex bone anatomies, making it an ideal test case for evaluating and refining fracture segmentation techniques.
For future iterations of the PENGWIN challenge, we aim to expand the challenge dataset and explore interactive segmentation approaches that integrate human decision-making with task-relevant information, thereby enhancing model performance in handling uncertainty and improving clinical applicability.